# D. Leykekhman, B. Vexler and J. Wagner

### v1

@@ Abstract
In this paper, we consider a state constrained optimal control problem governed by the transient Stokes equations. The state constraint is given by an $L^{2}$ functional in space, which is required to fulfill a pointwise bound in time. The discretization scheme for the Stokes equations consists of inf-sup stable finite elements in space and a discontinuous Galerkin method in time, for which we have recently established best approximation type error estimates. Using these error estimates, for the discrete control problem we establish error estimates and as a by-product we show an improved regularity for the optimal control. We complement our theoretical analysis with numerical results.
@@
keyword1; Keyword2; Keyword3; Keyword4.
@@ 1 Introduction
Here we assume that $\Omega\subset\mathbb{R}^{d}$, $d=2,3$, is a convex polygonal or polyhedral domain and $I=(0,T]$ is a bounded time interval. The control constraints are given by the constant vectors ${\bf{q}}_{a},{\bf{q}}_{b}\in{(\mathbb{R}\cup\{\pm\infty\})^{d}}$ and satisfy ${\bf{q}}_{a}<{\bf{q}}_{b}$. The constant scalar $\beta$ satisfies $\beta>0$, ${\bf w}(x)$ is a given function in $L^{2}(\Omega)^{d}$ and $\alpha>0$ is some regularization parameter. Note that for ease of presentation, we consider an optimal control problem with homogeneous initial data ${\bf{u}}(0)={\bf{0}}$, while all results also extend to the inhomogeneous case ${\bf{u}}(0)={\bf{u}}_{0}$. The main result of this paper states, that the error between the optimal control $\bar{{\bf{q}}}$ for the continuous problem and the optimal solution $\bar{{\bf{q}}}_{\sigma}$ of the discretized problem satisfies
$\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{\sigma}\rVert_{L^{2}(I;L^{2}(\Omega))}\leq\frac{C}{\sqrt{\alpha}}\,\ln{\frac{T}{k}}(k^{\frac{1}{2}}+h),$
@@ 2 Notation and Preliminary results
Throughout the paper, vector valued quantities and spaces will be indicated by boldface letters. We denote for a Banach space $X$ and $1\leq p\leq\infty$, by $L^{p}(I;X)$ the Bochner space of $X$-valued functions over $I$, whose $X$-norm is $p$-integrable w.r.t time. If $X$ is reflexive and $1\leq p<\infty$, there holds the following isomorphism, see [20, Corollary 1.3.22]
@@
$\mathbf{W}$ $:=L^{2}(I;\mathbf{V})\cap H^{1}(I;\mathbf{V}^{*})\hookrightarrow C(\bar{I};{\mathbf{V}_{0}}),$ (2.2)
@@
as it holds $\|A^{\frac{1}{2}}{\bf{u}}\|_{L^{2}(\Omega)}^{2}=\left(A{\bf{u}},{\bf{u}}\right)_{\Omega}=\left(\nabla{\bf{u}},\nabla{\bf{u}}\right)_{\Omega}=\|\nabla{\bf{u}}\|_{L^{2}(\Omega)}^{2}$. For the proof of $D(A^{\frac{1}{2}})=\mathbf{V}$, see [31, Ch. III, Lemma 2.2.1]. By its definition, we can extend $A$ to an operator (denoted by the same symbol) $A\colon\mathbf{V}\to\mathbf{V}^{*}$ yielding another isometric isomporphism between those spaces. Lastly, as
@@
$\|{\bf{u}}\|_{L^{2}(I;H^{2}(\Omega))}+\|{\bf{u}}\|_{L^{\infty}(I;\mathbf{V})}+\|\partial_{t}{\bf{u}}\|_{L^{2}(I;L^{2}(\Omega))}\leq C\|{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}.$
@@
The first part of this theorem is proven in [32, Chapter III, Theorem 1.1] and the remark on page 179 therein. The $H^{2}$ regularity part can be shown as in [32, Chapter III, Proposition 1.2] using the $H^{2}$ regularity result for the stationary Stokes equations in convex polygonal/polyhedral domains, proven in [9, Theorem 5.5, Theorem 6.3] see also [23, Theorem 2]. ∎
@@
see [3, Proposition 2.6], holding on fairly general domains, e.g. Lipschitz domains. If $\Omega$ is convex, the $H^{2}$ regularity results further imply that ${\bf{u}}\in L^{p}(I;\mathbf{V}_{2})$. Most often maximal parabolic regularity is treated in $L^{2}$ or $L^{q}$ setting in space, but it can also be extended to settings of weaker spacial regularity. The property (2.3) combined with [2, Lemma 11.4] yields that maximal parabolic regularity also holds in $\mathbf{V}^{*}$, i.e.
@@
If the right hand side ${\bf{q}}$ is regular enough, there exists an associated pressure to the weak solution of (2.4), formulated in divergence free spaces. The regularity of the pressure depends on the regularity of the right hand side and the velocity component of the solution. There holds the following result, see [3, Theorem 2.10, Corollary 2.11].
@@ 3 Continuous Optimal Control Problem
The first two claims are the direct consequences of Theorem 2.1. Note, that the regularity $\partial_{t}{\bf{u}}\in L^{2}(I;\mathbf{V}^{*})$ can only be obtained by bootstrapping if the right hand side is in $L^{2}(I;\mathbf{V}^{*})$. The third claim is obtained by using the maximal parabolic regularity and $H^{2}$ regularity of the stationary Stokes problem on convex polygonal/polyhedral domains, see [9, Theorem 5.5, Theorem 6.3]. ∎
@@
The operator $S$ is linear and coincides with its Frechét derivative, due to our choice to work with homogeneous inital data in the state equation (1.1b). The results presented in this work however also hold true in the inhomogeneous initial data case. In this case, the assumption $\beta>0$ then has to be generalized to $\beta>\left({\bf{u}}_{0},{\bf{w}}\right)_{\Omega}$.
Let us next give a characterization of the adjoint operator $S^{*}$. Theorem 2.1, together with [5, Proposition V.1.3], yields that for any ${\bf{q}}\in{\mathbf{Y}}$, there exists a unique solution ${\bf{u}}\in\mathbf{X}$ such that
@@
and let $\{{\bf{q}}_{n}\}$ and $\{{\bf{u}}_{n}:={\bf{u}}({\bf{q}}_{n})\}$ denote sequences of feasible controls with associated states, such that $J({\bf{q}}_{n},{\bf{u}}_{n})\to\mathcal{J}$ as $n\to\infty$. As it holds $J({\bf{q}}_{n},{\bf{u}}_{n})\leq{\mathcal{J}+1}$ for large enough $n$, there holds a bound $\|{\bf{q}}_{n}\|_{L^{2}(I;L^{2}(\Omega))}\leq C$ for all $n$. From Theorem 3.1, we obtain $\|{\bf{u}}_{n}\|_{L^{2}(I;H^{2}(\Omega))}+\|{\bf{u}}_{n}\|_{H^{1}(I;L^{2}(\Omega))}\leq C$. We can thus take a subsequence, denoted by the same index, such that
@@
These allow us to pass to the limit in the weak form of the state equation, showing that $\hat{\bf{u}}={\bf{u}}(\hat{\bf{q}})$. Furthermore, as $L^{2}(I;H^{2}(\Omega)^{d})\cap H^{1}(I;L^{2}(\Omega)^{d})\hookrightarrow C({\bar{I}};H^{s}(\Omega)^{d})$ compactly, for $s<1$, see [30, Corollary 8], by taking another subsequence, we obtain ${\bf{u}}_{n}\to\hat{\bf{u}}$ in $C({\bar{I}};L^{2}(\Omega)^{d})$. As $G_{\bf{w}}({\bf{u}}_{n})\leq\beta$ for all $t\in I$, this shows $G_{\bf{w}}(\hat{\bf{u}})\leq\beta$ for all $t\in I$. Lastly, it holds due to the lower semicontinuity of the norms
@@
We derive the optimality conditions from the reduced problem (3.5). By the generalized KKT condition (cf. [6, Theorem 5.2]) and the Slater condition from Assumption 1, i.e., the existence of $\tilde{{\bf{q}}}\in\mathbf{Q}_{ad}$ such that $\mathcal{G}(\tilde{{\bf{q}}})\in\operatorname{int}\mathcal{K}$, the optimality of $\bar{{\bf{q}}}$ is equivalent to the existence of a Lagrange multiplier $\bar{\mu}\in C(\bar{I})^{*}$ and the adjoint state
@@
$\left\langle\xi,\mu\right\rangle_{I\times\Omega}=\int_{I}\xi(t)\ \mathrm{d}\tilde{\mu}(t).$
@@
$\left\langle\xi,\mu\right\rangle_{I\times\Omega}=-\int_{I}\tilde{\mu}(t)\ \mathrm{d}\xi(t)+\tilde{\mu}(T)\xi(T)-\tilde{\mu}(0)\xi(0)=-\int_{I}\tilde{\mu}(t)\partial_{t}\xi(t)\ \mathrm{d}t-\tilde{\mu}(0)\xi(0)$
@@
The available regularity of $\bar{\bf{z}}$ immediately gives the claimed $L^{s}$ regularity in time of $\bar{\bf{q}}$. Note that due to the application of $P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}$, $\bar{\bf{q}}$ is in general not divergence free anymore and exhibits at most $W^{1,\infty}(\Omega)$ regularity in space, even if $\bar{\bf{z}}$ is smoother. We now turn towards showing continuity in time of $\bar{\bf{q}}$. Let us denote by []delimited-[][$\cdot$][ ⋅ ] the jump function w.r.t. time, i.e. $[\varphi](t)=\varphi(t+)-\varphi(t-)$. Since functions in $\mathrm{BV}(I;X)$ for any Banach space $X$, possess well-defined onesided limits, see [19, Propositions 2.1 & 2.2], we obtain from (3.10) that ${\bf{z}}_{\alpha}(t\pm)\in\mathbf{V}_{0}$ is well defined for any $t\in I$. Hence, by continuity of $P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}\colon L^{2}(\Omega)^{d}\to L^{2}(\Omega)^{d}$ we also have $[\bar{\bf{q}}]=[P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}({\bf{z}}_{\alpha})]\in L^{2}(\Omega)^{d}$ for all $t\in I$. By distinguishing the different cases, it is straightforward to verify, that for any ${\bf{v}}\in C(\bar{I};L^{2}(\Omega)^{d})+\mathrm{BV}(I;L^{2}(\Omega)^{d})$ it holds
@@
$\left(\bar{\bf{q}},{\bf{w}}\right)_{\Omega}+\left(\nabla\bar{\bf{u}},\nabla{\bf{w}}\right)_{\Omega}=\left(\partial_{t}\bar{\bf{u}},{\bf{w}}\right)_{\Omega}=\partial_{t}G_{\bf{w}}(\bar{\bf{u}})\quad\text{a.e. in }I.$
@@
where the last embedding holds due to [30, Corollary 8]. As further ${\bf{w}}\in\mathbf{V}_{2}\hookrightarrow C(\bar{\Omega})$, for any $t^{*}\in I$ it thus holds $\bar{\bf{z}}(t^{*}+),\bar{\bf{z}}(t^{*}-)\in C(\bar{\Omega})$. As $\bar{\bf{z}}(t^{*}+)(x^{*})=\bar{\bf{z}}(t^{*}-)(x^{*})={\bf{0}}$, and ${\bf{q}}_{a}<{\bf{0}}<{\bf{q}}_{b}$, there exists $\delta>0$, such that

### v2

@@ Abstract
In this paper, we consider a state constrained optimal control problem governed by the transient Stokes equations. The state constraint is given by an $L^{2}$ functional in space, which is required to fulfill a pointwise bound in time. The discretization scheme for the Stokes equations consists of inf-sup stable finite elements in space and a discontinuous Galerkin method in time, for which we have recently established best approximation type error estimates. Using these error estimates, for the discrete control problem we derive error estimates and as a by-product we show an improved regularity for the optimal control. We complement our theoretical analysis with numerical results.
@@
Stokes equations, instationary, optimal control, state constraints, error estimates
@@ 1 Introduction
Here we assume that $\Omega\subset\mathbb{R}^{d}$, $d=2,3$, is a convex polygonal or polyhedral domain and $I=(0,T]$ is a bounded time interval. In the objective function, ${\bf{u}}_{d}\in L^{2}(I\times\Omega)^{d}$ represents the desired state and $\alpha>0$ is the regularization parameter. The control constraints are given by the constant vectors ${\bf{q}}_{a},{\bf{q}}_{b}\in{(\mathbb{R}\cup\{\pm\infty\})^{d}}$ and satisfy ${\bf{q}}_{a}<{\bf{q}}_{b}$. In the state constraint, the constant scalar $\beta$ satisfies $\beta>0$ and ${\bf w}(x)$ is a given function in $L^{2}(\Omega)^{d}$. Note that for ease of presentation, we consider an optimal control problem with homogeneous initial data ${\bf{u}}(0)={\bf{0}}$, while all results also extend to the inhomogeneous case ${\bf{u}}(0)={\bf{u}}_{0}$. The main result of this paper states, that the error between the optimal control $\bar{{\bf{q}}}$ for the continuous problem and the optimal solution $\bar{{\bf{q}}}_{\sigma}$ of the discretized problem satisfies
$\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{\sigma}\rVert_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq\frac{C}{\sqrt{\alpha}}\,\ln{\frac{T}{k}}(k^{\frac{1}{2}}+h),$
@@ 2 Notation and Preliminary results
Throughout the paper, vector valued quantities and spaces will be indicated by boldface letters. We denote for a Banach space $X$ and $1\leq p\leq\infty$, by $L^{p}(I;X)$ the Bochner space of $X$-valued functions over $I$, whose $X$-norm is $p$-integrable w.r.t time. If $X$ is reflexive and $1\leq p<\infty$, there holds the following isomorphism, see [23, Corollary 1.3.22]
@@
$\mathbf{W}$ $:=L^{2}(I;\mathbf{V})\cap H^{1}(I;\mathbf{V}^{*})\hookrightarrow C(\bar{I};{\mathbf{V}_{0}}),$
@@
as it holds $\|A^{\frac{1}{2}}{\bf{u}}\|_{L^{2}(\Omega)}^{2}=\left(A{\bf{u}},{\bf{u}}\right)_{\Omega}=\left(\nabla{\bf{u}},\nabla{\bf{u}}\right)_{\Omega}=\|\nabla{\bf{u}}\|_{L^{2}(\Omega)}^{2}$. For the proof of $D(A^{\frac{1}{2}})=\mathbf{V}$, see [35, Ch. III, Lemma 2.2.1]. By its definition, we can extend $A$ to an operator (denoted by the same symbol) $A\colon\mathbf{V}\to\mathbf{V}^{*}$ yielding another isometric isomporphism between those spaces. Lastly, as
@@
$\|{\bf{u}}\|_{L^{2}(I;H^{2}(\Omega))}+\|{\bf{u}}\|_{L^{\infty}(I;\mathbf{V})}+\|\partial_{t}{\bf{u}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C\|{\bf{q}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}.$
@@
The first part of this theorem is proven in [36, Chapter III, Theorem 1.1] and the remark on page 179 therein. The $H^{2}$ regularity part can be shown as in [36, Chapter III, Proposition 1.2] using the $H^{2}$ regularity result for the stationary Stokes equations in convex polygonal/polyhedral domains, proven in [10, Theorem 5.5, Theorem 6.3] see also [26, Theorem 2]. ∎
@@
see [4, Proposition 2.6], holding on fairly general domains, e.g. Lipschitz domains. If $\Omega$ is convex, the $H^{2}$ regularity results further imply that ${\bf{u}}\in L^{p}(I;\mathbf{V}_{2})$. Most often maximal parabolic regularity is treated in $L^{2}$ or $L^{q}$ setting in space, but it can also be extended to settings of weaker spacial regularity. The property (2.2) combined with [3, Lemma 11.4] yields that maximal parabolic regularity also holds in $\mathbf{V}^{*}$, i.e.
@@
If the right hand side ${\bf{q}}$ is regular enough, there exists an associated pressure to the weak solution of (2.3), formulated in divergence free spaces. The regularity of the pressure depends on the regularity of the right hand side and the velocity component of the solution. There holds the following result, see [4, Theorem 2.10, Corollary 2.11].
@@ 3 Continuous Optimal Control Problem
The first two claims are the direct consequences of Theorem 2.1. Note, that the regularity $\partial_{t}{\bf{u}}\in L^{2}(I;\mathbf{V}^{*})$ can only be obtained by bootstrapping if the right hand side is in $L^{2}(I;\mathbf{V}^{*})$. The third claim is obtained by using the maximal parabolic regularity and $H^{2}$ regularity of the stationary Stokes problem on convex polygonal/polyhedral domains, see [10, Theorem 5.5, Theorem 6.3]. ∎
@@
The operator $S$ is linear and coincides with its Frechét derivative, due to our choice to work with homogeneous inital data in the state equation (1.1b). The results presented in this work however also hold true in the inhomogeneous initial data case. In this case, the assumption $\beta>0$ , on the parameter $\beta$ of the state constraint, then has to be generalized to $\beta>\left({\bf{u}}_{0},{\bf{w}}\right)_{\Omega}$.
Let us next give a characterization of the adjoint operator $S^{*}$. Theorem 2.1, together with [6, Proposition V.1.3], yields that for any ${\bf{q}}\in{\mathbf{Y}}$, there exists a unique solution ${\bf{u}}\in\mathbf{X}$ such that
@@
and let $\{{\bf{q}}_{n}\}$ and $\{{\bf{u}}_{n}:={\bf{u}}({\bf{q}}_{n})\}$ denote sequences of feasible controls with associated states, such that $J({\bf{q}}_{n},{\bf{u}}_{n})\to\mathcal{J}$ as $n\to\infty$. As it holds $J({\bf{q}}_{n},{\bf{u}}_{n})\leq{\mathcal{J}+1}$ for large enough $n$, there holds a bound $\|{\bf{q}}_{n}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C$ for all $n$. From Theorem 3.1, we obtain $\|{\bf{u}}_{n}\|_{L^{2}(I;H^{2}(\Omega))}+\|{\bf{u}}_{n}\|_{H^{1}(I;L^{2}(\Omega))}\leq C$. We can thus take a subsequence, denoted by the same index, such that
@@
These allow us to pass to the limit in the weak form of the state equation, showing that $\hat{\bf{u}}={\bf{u}}(\hat{\bf{q}})$. Furthermore, as $L^{2}(I;H^{2}(\Omega)^{d})\cap H^{1}(I;L^{2}(\Omega)^{d})\hookrightarrow C({\bar{I}};H^{s}(\Omega)^{d})$ compactly, for $s<1$, see [34, Corollary 8], by taking another subsequence, we obtain ${\bf{u}}_{n}\to\hat{\bf{u}}$ in $C({\bar{I}};L^{2}(\Omega)^{d})$. As $G_{\bf{w}}({\bf{u}}_{n})\leq\beta$ for all $t\in I$, this shows $G_{\bf{w}}(\hat{\bf{u}})\leq\beta$ for all $t\in I$. Lastly, it holds due to the lower semicontinuity of the norms
@@
We derive the optimality conditions from the reduced problem (3.5). By the generalized KKT condition (cf. [7, Theorem 5.2]) and the Slater condition from Assumption 1, i.e., the existence of $\tilde{{\bf{q}}}\in\mathbf{Q}_{ad}$ such that $\mathcal{G}(\tilde{{\bf{q}}})\in\operatorname{int}\mathcal{K}$, the optimality of $\bar{{\bf{q}}}$ is equivalent to the existence of a Lagrange multiplier $\bar{\mu}\in C(\bar{I})^{*}$ and the adjoint state
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left\langle\xi,\mu\right\rangle_{I}}=\int_{I}\xi(t)\ \mathrm{d}\tilde{\mu}(t).$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left\langle\xi,\mu\right\rangle_{I}}=-\int_{I}\tilde{\mu}(t)\ \mathrm{d}\xi(t)+\tilde{\mu}(T)\xi(T)-\tilde{\mu}(0)\xi(0)=-\int_{I}\tilde{\mu}(t)\partial_{t}\xi(t)\ \mathrm{d}t-\tilde{\mu}(0)\xi(0)$
@@
The available regularity of $\bar{\bf{z}}$ immediately gives the claimed $L^{s}$ regularity in time of $\bar{\bf{q}}$. Note that due to the application of $P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}$, $\bar{\bf{q}}$ is in general not divergence free anymore and exhibits at most $W^{1,\infty}(\Omega)$ regularity in space, even if $\bar{\bf{z}}$ is smoother. We now turn towards showing continuity in time of $\bar{\bf{q}}$. Let us denote by []delimited-[][$\cdot$][ ⋅ ] the jump function w.r.t. time, i.e. $[\varphi](t)=\varphi(t+)-\varphi(t-)$. Since functions in $\mathrm{BV}(I;X)$ for any Banach space $X$, possess well-defined onesided limits, see [20, Propositions 2.1 & 2.2], we obtain from (3.10) that ${\bf{z}}_{\alpha}(t\pm)\in\mathbf{V}_{0}$ is well defined for any $t\in I$. Hence, by continuity of $P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}\colon L^{2}(\Omega)^{d}\to L^{2}(\Omega)^{d}$ we also have $[\bar{\bf{q}}]=[P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}({\bf{z}}_{\alpha})]\in L^{2}(\Omega)^{d}$ for all $t\in I$. By distinguishing the different cases, it is straightforward to verify, that for any ${\bf{v}}\in C(\bar{I};L^{2}(\Omega)^{d})+\mathrm{BV}(I;L^{2}(\Omega)^{d})$ it holds
@@
$\left(\bar{\bf{q}},{\bf{w}}\right)_{\Omega}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\left(\nabla\bar{\bf{u}},\nabla{\bf{w}}\right)_{\Omega}=\left(\partial_{t}\bar{\bf{u}},{\bf{w}}\right)_{\Omega}=\partial_{t}G_{\bf{w}}(\bar{\bf{u}})\quad\text{a.e. in }I.$
@@
where the last embedding holds due to [34, Corollary 8]. As further ${\bf{w}}\in\mathbf{V}_{2}\hookrightarrow C(\bar{\Omega})$, for any $t^{*}\in I$ it thus holds $\bar{\bf{z}}(t^{*}+),\bar{\bf{z}}(t^{*}-)\in C(\bar{\Omega})$. As $\bar{\bf{z}}(t^{*}+)(x^{*})=\bar{\bf{z}}(t^{*}-)(x^{*})={\bf{0}}$, and ${\bf{q}}_{a}<{\bf{0}}<{\bf{q}}_{b}$, there exists $\delta>0$, such that